Polynomial Inequalities:

9) 2x^2 > 5x + 3

show number line and answer in interval notation


1
Expert's answer
2021-09-22T22:54:48-0400

Let us solve the equality "2x^2 > 5x + 3," which is equivalent to "2x^2 - 5x - 3>0," and hence to "(x-3)(2x+1)>0." The equation "(x-3)(2x+1)=0" has two roots "x_1=3,\ x_2=-\frac{1}2=-0.5."

Let us use number line test. For this draw the following intervals:



Since "2(-1)^2-5(-1)-3=4>0,\ 2\cdot0^2-5\cdot 0-3=-3<0," and "2\cdot 4^2-5\cdot 4-3=9>0," we conclude that the solutions of the inequality belongs to the real interval "(-\infty, -0.5)\cup(3,+\infty)."